# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ 7. Proof of the main results
Let $1<p<Q$ be fixed and for $\varepsilon\in(0,1]$ let us consider $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the minimizer of (1.10) in the set (1.11) which is the unique finite energy solution to (1.9a). As already recalled in the introduction, since $\Omega$ has $C^{2}$-smooth boundary by the interior and boundary regularity theory for $p$-harmonic function we have $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ (the gradient $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ being actually Hölder continuous up to the boundary) and it is strictly positive in $\overline{\Omega}$ by the strong maximum principle.
Next, we are going to prove that for $\varepsilon\in(0,1]$ when the domain $\Omega$ satisfies the uniform gauge-ball condition $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ each solution $v^{\varepsilon}_{p}$ satisfies (1.12) for a constant $C>0$ depending on $R_{0}$ and $Q$. To do this, first notice that the inequality holds on $\partial\Omega$ in view of Proposition 6.2 with some constant $\widetilde{C}>0$ depending only on $R_{0}$ and $Q$. Moreover, in view of Proposition 6.1 we have $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}=o(v^{\varepsilon}_{p})$ as $\|g\|\to\infty$, therefore we can apply Proposition 6.3 and the inequality (1.12) follows just by taking $C=\sqrt{2}\widetilde{C}$.
Finally, if $\varepsilon=0$ and $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ is the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) then the interior regularity theory for $p$-harmonic function in the sub-Riemannian Heisenberg group proved in [37] yields $v^{0}_{p}\in C(\Omega)$ and $\nabla_{0}v^{0}_{p}\in C(\Omega)$. Moreover $v^{\varepsilon}_{p}\to v^{0}_{p}$ as $\varepsilon\to 0$ strongly in $\dot{W}^{1,p}_{1,0}(\Omega)$ and $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\to|\nabla_{0}v^{\varepsilon}_{p}|_{0}$ a.e. in $\Omega$ because of Proposition 7.1. Since $d_{0}\leq d_{\varepsilon}$ for any $\varepsilon>0$, the functions $\{v^{\varepsilon}_{p}\}_{\varepsilon\in(0,1]}$ are $d_{0}$-equiLipschitz in $\mathbb{H}^{n}$ due to the bound (1.12) and the fact that $0<v^{\varepsilon}_{p}\leq 1$ on $\overline{\Omega}$. As a consequence $v^{\varepsilon}_{p}\to v^{0}_{p}$ locally uniformly on $\mathbb{H}^{n}$ by the Ascoli-Arzelà Theorem, hence $v^{0}_{p}\in C(\overline{\Omega})$, and (1.12) follows for $v^{0}_{p}$ from the same inequality for each $v^{\varepsilon}_{p}$ as $\varepsilon\to 0$. ∎
In order to prove the main result of the paper, i.e., Theorem 1.4, we need three auxiliary results. The first one relates weak solutions of (1.9a) and (1.6a).
@@
Let $1<p<Q$, $\varepsilon\in[0,1]$ and $v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega)$ nonegative and nonconstant. If $v^{\varepsilon}_{p}$ is a weak solution of (1.9a), i.e., for test functions $\psi\in W^{1,p}_{\varepsilon}(\Omega)$ with compact support in $\Omega$, then $v^{\varepsilon}_{p}$ is continuous in $\Omega$, it is strictly positive and $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ is continuous in $\Omega$. As a consequence, if $v^{\varepsilon}_{p}\not\equiv 0$ then the function $u^{\varepsilon}_{p}:=(1-p)\log(v^{\varepsilon}_{p})$ is well-defined, it is $d_{\varepsilon}$-locally Lipschitz and it is a weak solution of (1.6a), i.e., for test functions $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support in $\Omega$.
@@
First, to simplify the notation we just write throghout the proof $v=v^{\varepsilon}_{p}$ and $u=u^{\varepsilon}_{p}$. Note that by regularity theory recalled in the introduction both $v$ and $\nabla_{\varepsilon}v$ are continuous in $\Omega$, moreover $v>0$ in $\overline{\Omega}$ by the strong maximum principle.
@@
$-\int_{\Omega}|\nabla_{\varepsilon}u|^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}=\int_{\Omega}|\nabla u_{\varepsilon}|_{\varepsilon}^{p}\varphi\,,$
@@
We give the prooof only in the case $(\varepsilon,p)\to(0,1)$, as the two others are similar and even simpler, adapting the argument in [34] for the Euclidean case (see also [15] for the case $\mathbb{H}^{1}$ when $\varepsilon=0$).
@@
Note that, according to [14, Theorem 2.15], there exists a sequence of smooth functions $(w_{k})_{k\in\mathbb{N}}\subset C^{\infty}(V)$ such that
@@
Concerning the left hand side of (7.8), chosing $w=u$ and combining (1.13), the uniform convergence $u_{p}^{\varepsilon}\to u$ and (i) we infer that
[…]
@@
On the other hand, note that $\nabla_{0}u_{p}^{\varepsilon}\to\nabla_{0}u$-weakly in $L^{\infty}(\mathbb{H}^{n})$ because (1.13) holds uniformly and $u^{\varepsilon}_{p}\to u$ locally uniformly, whence weak lower semicontinuity and Young’s inequality give
@@
Let us fix $\varepsilon\in[0,1]$ and for any $1<p<Q$ let $v^{\varepsilon}_{p}$ be the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ and let $u^{\varepsilon}_{p}$ the corresponding functions given by (1.8). According to Theorem 1.2 we know that (1.12) holds for $\{v^{\varepsilon}_{p}\}$ for a constant $C>0$ independent of $p$ as well as of $\varepsilon$. Hence, in view of Corollary 1.3 functions in $\{u^{\varepsilon}_{p}\}$ satisfy the uniform bound (1.13), so that $\{u^{\varepsilon}_{p}\}_{p\in(1,Q)}$ are equi-Lipschitz with respect to $d_{\varepsilon}$, strictly positive in $\Omega$ and vanishing identically on $\mathbb{H}^{n}\setminus\Omega$. Moreover, they are weak solutions to (1.6a) due to Lemma 7.2, so that they minimize the functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}$ defined in (1.7) with respect to competitors $w\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ such that $w-u^{\varepsilon}_{p}$ is compactly supported because of Lemma 7.3.
@@
(7.14) $\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\leq v^{0}_{p}(g)\leq\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{-\gamma}\quad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,,$
where the second inequality trivially holds in $\overline{\Omega}\cap B_{\bar{R}}(g_{0})$ because $v^{0}_{p}\leq 1$ in $\mathbb{H}^{n}$. Taking logarithm and multiplying by $01-p<$ 0, as $p\to 1$ inequality (1.14) follows in the form stated in the case $\varepsilon=0$.
As for the case $\varepsilon>0$, concerning the upper bound for each $u^{\varepsilon}=\lim u^{\varepsilon}_{p}$ we rely on Proposition 5.1. Thus, for $\varepsilon\in(0,1]$ and $1<p\leq 2$, since $\Omega\subset\mathbb{H}^{n}$ satisfies assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ and $B_{R_{0}}(g_{0})\subset\Omega^{c}$ then for each $\bar{R}>2>R_{0}$ such that $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ there exists $\widehat{C}>1$ depending only on $R_{0}$ and $\bar{R}$ such that for $v^{\varepsilon}_{p}$, $u^{\varepsilon}_{p}=(1-p)\log v^{\varepsilon}_{p}$, we have
(7.15) $v^{\varepsilon}_{p}(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,.$
@@
When dealing with the lower bound in (1.14) for $\varepsilon\in(0,1]$ we rely instead on Proposition 5.5 so that inequality (5.11) holds in $\overline{\Omega}$ for some $C_{0}>1$ depending only on $R_{0}$, $\bar{R}$, $\partial\Omega$. Taking logarithm of both sides for $\gamma=\frac{Q-p}{p-1}$ and $g\in\overline{\Omega}$ we obtain
@@
which completes the proof of the lower bound. Note that the previous argument still applies in the case of the joint limit $(\varepsilon,p)\to(0,1)$, hence inequality (5.11) holds no matter how $u=\lim_{\varepsilon,p}u^{\varepsilon}_{p}$ is constructed.

### v3

@@ 8. Proof of the main results
Let $1<p\leq 2$ be fixed and for $\varepsilon\in(0,1]$ let us consider $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the minimizer of (1.10) in the set (1.11) which is the unique finite energy solution to (1.9a). As already recalled in the introduction, since $\Omega$ has $C^{2}$-smooth boundary by the interior and boundary regularity theory for $p$-harmonic function we have $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ (the gradient $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ being actually Hölder continuous up to the boundary) and it is strictly positive in $\overline{\Omega}$ by the strong maximum principle.
Next, we are going to prove that for $\varepsilon\in(0,1]$ when the domain $\Omega$ satisfies the uniform gauge-ball condition $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ each solution $v^{\varepsilon}_{p}$ satisfies (1.12) for a constant $C>0$ depending on $R_{0}$ and $Q$. To do this, first notice that the inequality holds on $\partial\Omega$ in view of Proposition 7.1 with some constant $\widetilde{C}>0$ depending only on $R_{0}$ and $Q$. Moreover, in view of Proposition 8.1 we have $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}=o(v^{\varepsilon}_{p})$ as $\|g\|\to\infty$, therefore we can apply Theorem 7.4 and the inequality (1.12) follows just by taking $C=\sqrt{\overline{M}^{2}+5QM_{\ast}^{2}+4Q}$, with $\overline{M}$ and $M_{*}$ as defined in (7.10) and (7.9) respectively.
Finally, if $\varepsilon=0$ and $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ is the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) then the interior regularity theory for $p$-harmonic function in the sub-Riemannian Heisenberg group proved in [39] yields $v^{0}_{p}\in C(\Omega)$ and $\nabla_{0}v^{0}_{p}\in C(\Omega)$. Moreover $v^{\varepsilon}_{p}\to v^{0}_{p}$ as $\varepsilon\to 0$ strongly in $\dot{W}^{1,p}_{1,0}(\Omega)$ and $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}\to|\nabla_{0}v^{\varepsilon}_{p}|_{0}$ a.e. in $\Omega$ because of Proposition 8.2. Since $d_{\varepsilon}\leq d_{0}$ for any $\varepsilon>0$, the functions $\{v^{\varepsilon}_{p}\}_{\varepsilon\in(0,1]}$ are $d_{0}$-equiLipschitz in $\mathbb{H}^{n}$ due to the bound (1.12) and the fact that $0<v^{\varepsilon}_{p}\leq 1$ on $\overline{\Omega}$. As a consequence $v^{\varepsilon}_{p}\to v^{0}_{p}$ locally uniformly on $\mathbb{H}^{n}$ by the Ascoli-Arzelà Theorem, hence $v^{0}_{p}\in C(\overline{\Omega})$, and (1.12) follows for $v^{0}_{p}$ from the same inequality for each $v^{\varepsilon}_{p}$ as $\varepsilon\to 0$. ∎
In order to prove the main result of the paper, i.e., Theorem 1.4, we need three auxiliary results which are actually still valid in the range $1<p<Q$. The first one relates weak solutions of (1.9a) and (1.6a).
@@
Let $1<p<Q$, $\varepsilon\in[0,1]$ and $v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega)$ nonnegative and nonconstant. If $v^{\varepsilon}_{p}$ is a weak solution of (1.9a), i.e., for test functions $\psi\in W^{1,p}_{\varepsilon}(\Omega)$ with compact support in $\Omega$, then $v^{\varepsilon}_{p}$ is continuous in $\Omega$, it is strictly positive and $\nabla_{\varepsilon}v^{\varepsilon}_{p}$ is continuous in $\Omega$. As a consequence, if $v^{\varepsilon}_{p}\not\equiv 0$ then the function $u^{\varepsilon}_{p}:=(1-p)\log(v^{\varepsilon}_{p})$ is well-defined, it is $d_{\varepsilon}$-locally Lipschitz and it is a weak solution of (1.6a), i.e., for test functions $\varphi\in W^{1,p}_{\varepsilon}(\Omega)\cap L^{\infty}(\Omega)$ with compact support in $\Omega$.
@@
First, to simplify the notation we just write throughout the proof $v=v^{\varepsilon}_{p}$ and $u=u^{\varepsilon}_{p}$. Note that by regularity theory recalled in the introduction both $v$ and $\nabla_{\varepsilon}v$ are continuous in $\Omega$, moreover $v>0$ in $\overline{\Omega}$ by the strong maximum principle.
@@
$\int_{\Omega}|\nabla_{\varepsilon}u|^{p-2}\langle\nabla_{\varepsilon}u,\nabla_{\varepsilon}\varphi\rangle_{\varepsilon}=\int_{\Omega}|\nabla u_{\varepsilon}|_{\varepsilon}^{p}\varphi\,$,
@@
We give the proof only in the case $(\varepsilon,p)\to(0,1)$, as the two others are similar and even simpler, adapting the argument in [36] for the Euclidean case (see also [16] for the case $\mathbb{H}^{1}$ when $\varepsilon=0$).
@@
Note that, according to [15, Theorem 2.15], there exists a sequence of smooth functions $(w_{k})_{k\in\mathbb{N}}\subset C^{\infty}(V)$ such that
@@
Concerning the left hand side of (8.11), choosing $w=u$ and combining (1.13), the uniform convergence $u_{p}^{\varepsilon}\to u$ and (i) we infer that
$\limsup_{k\to\infty}\limsup_{(\varepsilon,p)\to(0,1)}\int_{\rm{supp}(\eta)}\eta|u_{p}^{\varepsilon}-w_{k}||\nabla_{\varepsilon}u_{p}^{\varepsilon}|_{\varepsilon}^{p}\leq\limsup_{k\to\infty}\int_{\rm{supp}(\eta)}C\eta|u-w_{k}|=0\,$,
@@
On the other hand, note that $\nabla_{0}u_{p}^{\varepsilon}\to\nabla_{0}u$ *-weakly in $L^{\infty}(\mathbb{H}^{n})$ because (1.13) holds uniformly and $u^{\varepsilon}_{p}\to u$ locally uniformly, whence weak lower semicontinuity and Young’s inequality give
@@
Let us fix $\varepsilon\in[0,1]$ and for any $1<p\leq 2$ let $v^{\varepsilon}_{p}$ be the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ and let $u^{\varepsilon}_{p}$ the corresponding functions given by (1.8). According to Theorem 1.2 we know that (1.12) holds for $\{v^{\varepsilon}_{p}\}$ for a constant $C>0$ independent of $p$ as well as of $\varepsilon$. Hence, in view of Corollary 1.3 functions in $\{u^{\varepsilon}_{p}\}$ satisfy the uniform bound (1.13), so that $\{u^{\varepsilon}_{p}\}_{p\in(1,Q)}$ are equi-Lipschitz with respect to $d_{\varepsilon}$, strictly positive in $\Omega$ and vanishing identically on $\mathbb{H}^{n}\setminus\Omega$. Moreover, they are weak solutions to (1.6a) due to Lemma 8.3, so that they minimize the functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}$ defined in (1.7) with respect to competitors $w\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ such that $w-u^{\varepsilon}_{p}$ is compactly supported because of Lemma 8.4.
@@
(8.17) $\left(\frac{\|g_{0}^{-1}*g\|}{R_{\ast}}\right)^{-\gamma}\leq v^{0}_{p}(g)\leq\left(\frac{\|g_{0}^{-1}*g\|}{\bar{R}}\right)^{-\gamma}\quad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,$,
where the second inequality trivially holds in $\overline{\Omega}\cap B_{\bar{R}}(g_{0})$ because $v^{0}_{p}\leq 1$ in $\mathbb{H}^{n}$. Taking logarithm and multiplying by 1p<01-p<0, as $p\to 1$ inequality (1.14) follows in the form stated in the case $\varepsilon=0$.
As for the case $\varepsilon>0$, concerning the upper bound for each $u^{\varepsilon}=\lim u^{\varepsilon}_{p}$ we rely on Proposition 5.1. Thus, for $\varepsilon\in(0,1]$ and $1<p\leq 2$, since $\Omega\subset\mathbb{H}^{n}$ satisfies assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ and for each $R_{\ast}<2<\bar{R}$ with $B_{R_{\ast}}(g_{0})\subset\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ there exists $\widehat{C}>1$ depending only on $Q$, $R_{\ast}$ and $\bar{R}$ such that for $v^{\varepsilon}_{p}$, $u^{\varepsilon}_{p}=(1-p)\log v^{\varepsilon}_{p}$, we have
(8.18) $v^{\varepsilon}_{p}(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{\ast}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,.$
@@
[…]
@@
which completes the proof of the lower bound. Note that the previous argument still applies in the case of the joint limit $(\varepsilon,p)\to(0,1)$, since inequality (5.11) can be applied no matter how $u=\lim_{\varepsilon,p}u^{\varepsilon}_{p}$ is constructed.
Acknowledgments
We thank Luca Capogna and Giovanna Citti for several interesting and useful discussions on the content of some results of Section 4 and the connections with their papers [9] and [10].
A.P. is partially funded by PRIN project 2022PJ9EFL “Geometric Measure Theory: Structure of Singular Measures, Regularity Theory and Applications in the Calculus of Variations”. E.V. is partially funded by PRIN project 2022R537CS $NO^{3}$ - Nodal Optimization, NOnlinear elliptic equations, NOnlocal geometric problems, with a focus on regularity”.
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.